# multimotion estimation problem MOT multiobject tracking

### v3

@@ Abstract
Visual motion estimation is a well-studied challenge in autonomous navigation. Recent work has focused on addressing multimotion estimation in highly dynamic environments. These environments not only comprise multiple, complex motions but also tend to exhibit significant occlusion. Estimating third-party motions simultaneously with the sensor egomotion is difficult because an object’s observed motion consists of both its true motion and the sensor motion. Most previous works in multimotion estimation simplify this problem by relying on appearance-based object detection or application-specific motion constraints. These approaches are effective in specific applications and environments but do not generalize well to the full multimotion estimation problem (MEP). This paper presents Multimotion Visual Odometry (MVO), a multimotion estimation pipeline that estimates the full $SE\left(3\right)$ trajectory of every motion in the scene, including the sensor egomotion, without relying on appearance-based information. MVO extends the traditional visual odometry (VO) pipeline with multimotion segmentation and tracking techniques. It uses physically founded motion priors to extrapolate motions through temporary occlusions and identify the reappearance of motions through motion closure. Evaluations on real-world data from the Oxford Multimotion Dataset (OMD) and the KITTI Vision Benchmark Suite demonstrate that MVO achieves good estimation accuracy compared to similar approaches and is applicable to a variety of multimotion estimation challenges.
@@ 2 Background
Dynamic environments consist of the static background, a moving observer (i.e., sensor), and one or more independent, third-party motions. The observed motions include both the objects’ real-world motion and the sensor egomotion, so the static scene appears to move with the inverse of the egomotion trajectory. Fully addressing this MEP requires both segmentation, i.e., clustering points according to their movement between observations, and estimation, i.e., calculating the motion of a cluster of points. The interdependence of these tasks creates a chicken-and-egg problem: the motion trajectories must be estimated from groups of points, but the segmentation of those groups depends on the available motion estimates. This interdependence is addressed in single-motion VO systems by using heuristics (e.g., number of features) to select the egomotion and ignore the other motions in the scene (Section 2.1). Appearance-based tracking techniques are often used to track multiple dynamic objects, but currently do not accurately estimate the underlying motion of generic objects (Section 2.2). These heuristic- and appearance-based techniques are not readily extensible to general multimotion estimation problems and analyzing multiple independently moving bodies remains a challenging problem for state-of-the-art vision systems (Section 2.3).
@@ 2.1 Egomotion Estimation
VO estimates the motion of a camera relative to its static environment (i.e., egomotion) and has a long history in robotics starting with the motion-estimation pipeline presented by Moravec (1980). The camera egomotion is calculated from a stream of images by assuming it corresponds to the dominant motion in the scene. The egomotion is estimated by first isolating regions of the scene that move according to the dominant motion and assuming they are static. Those regions are then used to estimate the egomotion trajectory relative to those static points and the rest of the scene is usually ignored as noise. Recent developments in VO focus on improving the static segmentation (Nistér, 2004; Sabzevari and Scaramuzza, 2016). The estimation can also be improved by using more advanced estimation techniques, such as dense or direct estimation methods (Valgaerts et al., 2012; Engel et al., 2017) and continuous $SE\left(3\right)$ motion priors (Anderson and Barfoot, 2015; Tang et al., 2019). These techniques make the egomotion estimation more robust and accurate, but few methods extend them to also analyze the dynamic parts of the scene that VO rejects.
@@ 2.2 Multiobject Tracking (MOT)
multiobject tracking (MOT) techniques are often used in tandem with VO to track multiple dynamic objects over time, usually through a tracking-by-detection paradigm (Milan et al., 2016). They use appearance-based object detectors to find instances of specific object classes in each frame. They then seek to accurately associate present and past detections to form simple $\mathbb{R}^{3}$ trajectories. These techniques are limited by the quality of their detectors, and are often constrained to tracking application-specific object classes. MOT algorithms often track objects through occlusions by modeling object interactions (Yang et al., 2011) or employing simple motion models to extrapolate object positions (Mitzel et al., 2010). These techniques are limited by the object representations used by the target detectors, and tend to be constrained to simple motion models that do not adequately represent the motion of an object. Fully addressing the MEP requires applying more expressive motion estimation techniques, such as those used to estimate egomotion, to the other third-party motions in a scene.
@@ 2.3 Multimotion Estimation
In contrast to MOT, multimotion estimation focuses on the motions in the scene rather than the objects that generate them. This involves both segmenting observed points into independent rigid objects and estimating the $SE\left(3\right)$ trajectories of those objects. Several multimotion estimation approaches have been developed using techniques such as scene flow clustering, simultaneous localization and mapping (SLAM) frameworks, and energy-minimization, but none has fully addressed the general MEP. Lenz et al. (2011) use sparse scene flow to detect and track multiple dynamic objects from a dynamic, vehicle-mounted camera. The approach operates on stereo image pairs and clusters sparse points based on their scene flow. These clusters are then used to track objects through $\mathbb{R}^{3}$ space. The approach is not limited to tracking specific classes of objects and is robust to some partial occlusions, but it requires objects to be of limited size and in contact with the ground plane. Scene flow also estimates pixel-wise translational motion, so estimating the motion of rotating objects requires further segmentation and estimation. Menze and Geiger (2015) instead model the world as piecewise-planar superpixels. The approach simultaneously calculates both scene flow and the homographies defining the motion of the planes in the scene, which are modeled as full $SE\left(3\right)$ motions. Similarly, Quiroga et al. (2014) and Jaimez et al. (2017) define optimization frameworks that estimate scene flow by modeling the underlying motions as full $SE\left(3\right)$ motions. This formulation is more robust to rotations but it complicates the MEP by introducing additional constraints to estimate the pixel-wise velocity of every point in the scene. Each of these techniques also relies on accurate RGB-D sensing and an independent egomotion estimation pipeline, which can lead to errors when large portions of the scene are moving. Wang et al. (2007) extend the traditional SLAM formulation to include MOT techniques in their SLAMMOT framework using lidar sensors and monocular and stereo cameras (Lin and Wang, 2010). The traditional SLAM state, consisting of the $SE\left(3\right)$ sensor pose and a static map of the environment, is extended to include the range, bearing, and linear velocity of any tracked objects in the scene. Like most 3D MOT algorithms, SLAMMOT only estimates the $\mathbb{R}^{3}$ position of third-party objects, rather than their full $SE\left(3\right)$ pose. CubeSLAM (Yang and Scherer, 2019) similarly extends the SLAM framework to track the full $SE\left(3\right)$ pose of 3D bounding boxes for each detected object, but not the pose of the object itself.
@@
Rather than tracking 3D bounding boxes, Huang et al. (2019) estimate bulk point motions in ClusterSLAM, a multibody dynamic SLAM backend that was extended to a full multimotion estimation pipeline in ClusterVO (Huang et al., 2020). The pipeline clusters tracked 3D points according to their bulk motion but relies on a semantic object detector to inform its data association and tracking. Visual Direct Object-aware SLAM (VDO-SLAM; Henein et al., 2020; Zhang et al., 2020a, b) also extends the SLAM formulation to estimate the full $SE\left(3\right)$ pose of other objects in the scene, but uses an external instance segmentation algorithm to segment the image into its constituent objects first. DynaSLAM II (Bescos et al., 2021) similarly uses front-end instance segmentation to guide feature matching and segmentation before estimating the $SE\left(3\right)$ trajectory of each tracked object. ClusterVO, Visual Direct Object-aware SLAM (VDO-SLAM), and DynaSLAM II are able to estimate the motion of each object more accurately than techniques that only track bounding boxes, but they are reliant on significant front-end preprocessing, which limits their applicability to the general MEP. In contrast, DymSLAM (Wang et al., 2020) is a SLAM framework without any significant front-end processing. It uses spectral clustering to segment dynamic motions and estimate their $SE\left(3\right)$ motions but primarily focuses on dense scene reconstruction which is separate from the MEP. Instead of building on the traditional SLAM framework, Isack and Boykov (2012) demonstrate an energy-based multimodel-fitting framework that can be used to segment motions in a frame-to-frame manner. Roussos et al. (2012) use this framework to simultaneously estimate depth maps and object motions from a monocular camera while also performing object tracking. The technique minimizes a complex cost function involving photometric consistency, geometric smoothness in depth, spatial smoothness in image space, and a minimum-description-length term that promotes a compact solution. The approach proceeds in an offline, full-batch manner and its additional focus on dense reconstruction, combined with its initialization requirements, make it ill-suited for MEP applications. Other techniques use RGB-D sensors to segment and estimate motions while also performing suitable tracking for object reconstruction. Rünz and Agapito (2017) use an RGB-D camera to segment and track targets while simultaneously fusing 3D object models. The technique combines motion segmentation with object instance segmentation, which relies on predefined class-based object detectors. Rünz et al. (2018) extend this work to real-time processing by improving the efficiency of the semantic segmentation, and Xu et al. (2019) define a similar system using a volumetric representation, rather than surface normals. These techniques represent significant progress in addressing the MEP, with the added ability to fuse 3D models of the tracked objects; but they are reliant on high-quality, dense depth sensing and they are limited in the number of active models they can reasonably process. Qiu et al. (2019) and Eckenhoff et al. (2019) address the MEP using monocular cameras. Monocular observations are underconstrained and a separate scale parameter must be estimated using the inertial measurement unit (IMU). This scale parameter is valid for the egomotion of the camera, but estimating the scale for third-party motions is difficult and leads to several degenerate cases, such as when the object and camera motions are colinear. The performance of these techniques is impressive given the limitations of the sensors, but they are not broadly applicable to the general MEP. In contrast to these approaches, MVO fully addresses the MEP in an online manner using only a rigid-motion assumption. The pipeline estimates the full $SE\left(3\right)$ trajectory of every motion in a complex, dynamic scene without any a priori information about object number, appearance, or motion. It also operates directly on sparse 3D points and is therefore applicable to a variety of 3D sensors. MVO can incorporate physically founded $SE\left(3\right)$ motion priors to extrapolate motions accurately through occlusions and achieves motion-based tracking through a motion closure procedure.
@@ 3 Simultaneous $SE\left(3\right)$ Estimation and Segmentation
Fully addressing the MEP requires both motion estimation and segmentation. Motion estimation involves calculating the motion of a set of points. Motion segmentation involves clustering points according to their movement. The interdependence of estimation and segmentation leads to a chicken-and-egg problem: the motion trajectories must be estimated from groups of points, but the segmentation of those groups depends on the available motion estimates. In single-motion VO systems this is addressed by using heuristics (e.g., number of features) to select the egomotion and ignore the other motions in the scene. These heuristics are not readily extensible to multimotion estimation problems and analyzing multiple independently moving bodies remains a challenging problem for state-of-the-art vision systems. MVO addresses the MEP by casting motion segmentation as a multilabeling problem where a label represents a motion trajectory. It iteratively segments and estimates motions using these labels in an alternating fashion until the segmentation converges. The set of motion labels adapts to the motions in the scene without making any a priori assumptions about object number, appearance, or motion. MVO does not use any assumptions or heuristics to identify the static background until after the segmentation converges, so each motion is estimated as if it corresponds to the true egomotion. After the segmentation converges, the true egomotion is selected and used to reestimate the true third-party motions in a geocentric frame (Section 4). The pipeline operates directly on tracked 3D observations, i.e., tracklets, from a 3D sensor. It iteratively generates labels and segments those observations according to their motion (Fig. 2). The tracklets are embedded in a graph structure that forms the basis of the motion segmentation (Section 3.2). New labels are proposed when the motion of tracklets assigned to a given label may be more accurately explained by multiple trajectories (Section 3.3). Motion labels are assigned to each tracklet by minimizing a cost functional that incorporates reprojection residual, graph smoothness, and model complexity (Section 3.4). Tracklets whose motions are not well explained by any other label are assigned to an outlier label and redundant and oversegmented labels are then merged (Section 3.4.4). Once the segmentation converges, the final labels are then sanitized and any remaining outliers are rejected (Section 3.5). Initial versions of these ideas were first presented in Judd et al. (2018a) and Judd et al. (2018b), and they are further developed here.

### v4

@@ Abstract
Visual motion estimation is a well-studied challenge in autonomous navigation. Recent work has focused on addressing multimotion estimation in highly dynamic environments. These environments not only comprise multiple, complex motions but also tend to exhibit significant occlusion.
Estimating third-party motions simultaneously with the sensor egomotion is difficult because an object’s observed motion consists of both its true motion and the sensor motion. Most previous works in multimotion estimation simplify this problem by relying on appearance-based object detection or application-specific motion constraints. These approaches are effective in specific applications and environments but do not generalize well to the full multimotion estimation problem (MEP).
This paper presents Multimotion Visual Odometry (MVO), a multimotion estimation pipeline that estimates the full $SE\left(3\right)$ trajectory of every motion in the scene, including the sensor egomotion, without relying on appearance-based information. MVO extends the traditional visual odometry (VO) pipeline with multimotion segmentation and tracking techniques. It uses physically founded motion priors to extrapolate motions through temporary occlusions and identify the reappearance of motions through motion closure. Evaluations on real-world data from the Oxford Multimotion Dataset (OMD) and the KITTI Vision Benchmark Suite demonstrate that MVO achieves good estimation accuracy compared to similar approaches and is applicable to a variety of multimotion estimation challenges.
@@ 1 Introduction
Third-party dynamic motions are more difficult to estimate than the sensor egomotion because their observed motions comprise both their arbitrary real-world motions and the sensor egomotion. This MEP is often simplified by constraining motions according to kinematic assumptions (e.g., Sabzevari and Scaramuzza, 2016), or by first isolating and estimating the sensor egomotion and then compensating for it to estimate the remaining third-party motions in the scene (e.g., Jaimez et al., 2017). These techniques can be successful in specific applications, but few generalized approaches have been proposed to address the full MEP.
[…]
The rest of the paper is organized as follows. Section 2 explores the details of the MEP and the current approaches for addressing it. Section 3 describes how the MVO pipeline extends traditional VO techniques to multimotion estimation. Section 4 demonstrates how different types of state estimators and motion priors can be used within the MVO pipeline. Section 5 discusses how to address the challenges involved in implementing MVO as a sliding-window estimator, particularly with respect to occlusions. Sections 6 and 7 evaluate and discuss the performance of the MVO pipeline quantitatively and qualitatively using the Oxford Multimotion Dataset (OMD; Judd and Gammell, 2019a) and the KITTI Vision Benchmark Suite (Geiger et al., 2012).
This paper is a continuation of ideas that were first presented at the Joint Industry and Robotics CDTs Symposium (Judd et al., 2018b) and the Long-term Human Motion Prediction Workshop (Judd and Gammell, 2019b) and were published in Judd et al. (2018a), Judd (2019), and Judd and Gammell (2020). It makes the following specific contributions:
@@ 2 Background
Dynamic environments consist of the static background, a moving observer (i.e., sensor), and one or more independent, third-party motions. The observed motions include both the objects’ real-world motion and the sensor egomotion, so the static scene appears to move with the inverse of the egomotion trajectory.
Fully addressing this MEP requires both segmentation, i.e., clustering points according to their movement between observations, and estimation, i.e., calculating the motion of a cluster of points. The interdependence of these tasks creates a chicken-and-egg problem: the motion trajectories must be estimated from groups of points, but the segmentation of those groups depends on the available motion estimates.
This interdependence is addressed in single-motion VO systems by using heuristics (e.g., number of features) to select the egomotion and ignore the other motions in the scene (Section 2.1). Appearance-based tracking techniques are often used to track multiple dynamic objects, but currently do not accurately estimate the underlying motion of generic objects (Section 2.2). These heuristic- and appearance-based techniques are not readily extensible to general multimotion estimation problems and analyzing multiple independently moving bodies remains a challenging problem for state-of-the-art vision systems (Section 2.3).
@@ 2.1 Egomotion Estimation
VO estimates the motion of a camera relative to its static environment (i.e., egomotion) and has a long history in robotics starting with the motion-estimation pipeline presented by Moravec (1980). The camera egomotion is calculated from a stream of images by assuming it corresponds to the dominant motion in the scene.
The egomotion is estimated by first isolating regions of the scene that move according to the dominant motion and assuming they are static. Those regions are then used to estimate the egomotion trajectory relative to those static points and the rest of the scene is usually ignored as noise.
Recent developments in VO focus on improving the static segmentation (Nistér, 2004; Sabzevari and Scaramuzza, 2016). The estimation can also be improved by using more advanced estimation techniques, such as dense or direct estimation methods (Valgaerts et al., 2012; Engel et al., 2017) and continuous $SE\left(3\right)$ motion priors (Anderson and Barfoot, 2015; Tang et al., 2019). These techniques make the egomotion estimation more robust and accurate, but few methods extend them to also analyze the dynamic parts of the scene that VO rejects.
@@ 2.2 Multiobject Tracking (MOT)
multiobject tracking (MOT) techniques are often used in tandem with VO to track multiple dynamic objects over time, usually through a tracking-by-detection paradigm (Milan et al., 2016). They use appearance-based object detectors to find instances of specific object classes in each frame. They then seek to accurately associate present and past detections to form simple $\mathbb{R}^{3}$ trajectories. These techniques are limited by the quality of their detectors, and are often constrained to tracking application-specific object classes.
MOT algorithms often track objects through occlusions by modeling object interactions (Yang et al., 2011) or employing simple motion models to extrapolate object positions (Mitzel et al., 2010). These techniques are limited by the object representations used by the target detectors, and tend to be constrained to simple motion models that do not adequately represent the motion of an object. Fully addressing the MEP requires applying more expressive motion estimation techniques, such as those used to estimate egomotion, to the other third-party motions in a scene.
@@ 2.3 Multimotion Estimation
In contrast to MOT, multimotion estimation focuses on the motions in the scene rather than the objects that generate them. This involves both segmenting observed points into independent rigid objects and estimating the $SE\left(3\right)$ trajectories of those objects. Several multimotion estimation approaches have been developed using techniques such as scene flow clustering, simultaneous localization and mapping (SLAM) frameworks, and energy-minimization, but none has fully addressed the general MEP.
Lenz et al. (2011) use sparse scene flow to detect and track multiple dynamic objects from a dynamic, vehicle-mounted camera. The approach operates on stereo image pairs and clusters sparse points based on their scene flow. These clusters are then used to track objects through $\mathbb{R}^{3}$ space. The approach is not limited to tracking specific classes of objects and is robust to some partial occlusions, but it requires objects to be of limited size and in contact with the ground plane. Scene flow also estimates pixel-wise translational motion, so estimating the motion of rotating objects requires further segmentation and estimation.
Menze and Geiger (2015) instead model the world as piecewise-planar superpixels. The approach simultaneously calculates both scene flow and the homographies defining the motion of the planes in the scene, which are modeled as full $SE\left(3\right)$ motions. Similarly, Quiroga et al. (2014) and Jaimez et al. (2017) define optimization frameworks that estimate scene flow by modeling the underlying motions as full $SE\left(3\right)$ motions. This formulation is more robust to rotations but it complicates the MEP by introducing additional constraints to estimate the pixel-wise velocity of every point in the scene. Each of these techniques also relies on accurate RGB-D sensing and an independent egomotion estimation pipeline, which can lead to errors when large portions of the scene are moving.
Wang et al. (2007) extend the traditional SLAM formulation to include MOT techniques in their SLAMMOT framework using lidar sensors and monocular and stereo cameras (Lin and Wang, 2010). The traditional SLAM state, consisting of the $SE\left(3\right)$ sensor pose and a static map of the environment, is extended to include the range, bearing, and linear velocity of any tracked objects in the scene. Like most 3D MOT algorithms, SLAMMOT only estimates the $\mathbb{R}^{3}$ position of third-party objects, rather than their full $SE\left(3\right)$ pose. CubeSLAM (Yang and Scherer, 2019) similarly extends the SLAM framework to track the full $SE\left(3\right)$ pose of 3D bounding boxes for each detected object, but not the pose of the object itself.
@@
Rather than tracking 3D bounding boxes, Huang et al. (2019) estimate bulk point motions in ClusterSLAM, a multibody dynamic SLAM backend that was extended to a full multimotion estimation pipeline in ClusterVO (Huang et al., 2020). The pipeline clusters tracked 3D points according to their bulk motion but relies on a semantic object detector to inform its data association and tracking. Visual Direct Object-aware SLAM (VDO-SLAM; Henein et al., 2020; Zhang et al., 2020a, b) also extends the SLAM formulation to estimate the full $SE\left(3\right)$ pose of other objects in the scene, but uses an external instance segmentation algorithm to segment the image into its constituent objects first. DynaSLAM II (Bescos et al., 2021) similarly uses front-end instance segmentation to guide feature matching and segmentation before estimating the $SE\left(3\right)$ trajectory of each tracked object. ClusterVO, Visual Direct Object-aware SLAM (VDO-SLAM), and DynaSLAM II are able to estimate the motion of each object more accurately than techniques that only track bounding boxes, but they are reliant on significant front-end preprocessing, which limits their applicability to the general MEP. In contrast, DymSLAM (Wang et al., 2020) is a SLAM framework without any significant front-end processing. It uses spectral clustering to segment dynamic motions and estimate their $SE\left(3\right)$ motions but primarily focuses on dense scene reconstruction which is separate from the MEP.
Instead of building on the traditional SLAM framework, Isack and Boykov (2012) demonstrate an energy-based multimodel-fitting framework that can be used to segment motions in a frame-to-frame manner. Roussos et al. (2012) use this framework to simultaneously estimate depth maps and object motions from a monocular camera while also performing object tracking. The technique minimizes a complex cost function involving photometric consistency, geometric smoothness in depth, spatial smoothness in image space, and a minimum-description-length term that promotes a compact solution. The approach proceeds in an offline, full-batch manner and its additional focus on dense reconstruction, combined with its initialization requirements, make it ill-suited for MEP applications.
Other techniques use RGB-D sensors to segment and estimate motions while also performing suitable tracking for object reconstruction. Rünz and Agapito (2017) use an RGB-D camera to segment and track targets while simultaneously fusing 3D object models. The technique combines motion segmentation with object instance segmentation, which relies on predefined class-based object detectors. Rünz et al. (2018) extend this work to real-time processing by improving the efficiency of the semantic segmentation, and Xu et al. (2019) define a similar system using a volumetric representation, rather than surface normals. These techniques represent significant progress in addressing the MEP, with the added ability to fuse 3D models of the tracked objects; but they are reliant on high-quality, dense depth sensing and they are limited in the number of active models they can reasonably process.
Qiu et al. (2019) and Eckenhoff et al. (2019) address the MEP using monocular cameras. Monocular observations are underconstrained and a separate scale parameter must be estimated using the inertial measurement unit (IMU). This scale parameter is valid for the egomotion of the camera, but estimating the scale for third-party motions is difficult and leads to several degenerate cases, such as when the object and camera motions are colinear. The performance of these techniques is impressive given the limitations of the sensors, but they are not broadly applicable to the general MEP.
In contrast to these approaches, MVO fully addresses the MEP in an online manner using only a rigid-motion assumption. The pipeline estimates the full $SE\left(3\right)$ trajectory of every motion in a complex, dynamic scene without any a priori information about object number, appearance, or motion. It also operates directly on sparse 3D points and is therefore applicable to a variety of 3D sensors. MVO can incorporate physically founded $SE\left(3\right)$ motion priors to extrapolate motions accurately through occlusions and achieves motion-based tracking through a motion closure procedure.
@@ 3 Simultaneous $SE\left(3\right)$ Estimation and Segmentation
Fully addressing the MEP requires both motion estimation and segmentation. Motion estimation involves calculating the motion of a set of points. Motion segmentation involves clustering points according to their movement.
The interdependence of estimation and segmentation leads to a chicken-and-egg problem: the motion trajectories must be estimated from groups of points, but the segmentation of those groups depends on the available motion estimates. In single-motion VO systems this is addressed by using heuristics (e.g., number of features) to select the egomotion and ignore the other motions in the scene. These heuristics are not readily extensible to multimotion estimation problems and analyzing multiple independently moving bodies remains a challenging problem for state-of-the-art vision systems.
MVO addresses the MEP by casting motion segmentation as a multilabeling problem where a label represents a motion trajectory. It iteratively segments and estimates motions using these labels in an alternating fashion until the segmentation converges. The set of motion labels adapts to the motions in the scene without making any a priori assumptions about object number, appearance, or motion.
MVO does not use any assumptions or heuristics to identify the static background until after the segmentation converges, so each motion is estimated as if it corresponds to the true egomotion. After the segmentation converges, the true egomotion is selected and used to reestimate the true third-party motions in a geocentric frame (Section 4).
The pipeline operates directly on tracked 3D observations, i.e., tracklets, from a 3D sensor. It iteratively generates labels and segments those observations according to their motion (Fig. 2). The tracklets are embedded in a graph structure that forms the basis of the motion segmentation (Section 3.2). New labels are proposed when the motion of tracklets assigned to a given label may be more accurately explained by multiple trajectories (Section 3.3). Motion labels are assigned to each tracklet by minimizing a cost functional that incorporates reprojection residual, graph smoothness, and model complexity (Section 3.4). Tracklets whose motions are not well explained by any other label are assigned to an outlier label and redundant and oversegmented labels are then merged (Section 3.4.4). Once the segmentation converges, the final labels are then sanitized and any remaining outliers are rejected (Section 3.5). Initial versions of these ideas were first presented in Judd et al. (2018a) and Judd et al. (2018b), and they are further developed here.
@@ 3.1 Notation
where $C_{k}$ refers to the observing sensor frame at time $k$ (Fig. 3, right), and $\boldsymbol{p}^{j_{k}C_{k}}_{C_{k}}\in\mathbb{R}^{4\times 1}$ is the homogeneous representation of the Euclidean vector, $\mathbf{p}^{j_{k}C_{k}}_{C_{k}}\in\mathbb{R}^{3\times 1}$.
@@
The set of tracklets visible in a scene, $\mathscr{P}$, comprises the subsets of tracklets assigned to each motion label, $\mathscr{P}_{\ell{}}$, i.e., the support for label $\ell{}$ The current set of all motion labels available for assignment is given by $\mathscr{L}$.